NicktheBlade16:

PLZ HELP, IT'S ALGEBRA 2: Daniel is looking over some data regarding the temperatures of countries and how the temperature relates to the number of mosquitoes. The data are represented by the exponential function f(x) = 2x, where x is the number of mosquitoes. Explain how he can convert this equation to a logarithmic function when the temperature is 64 degrees.

Astrid1:

Okay to start this off we can do : \[f(x)=2^x\] We're going to turn f(x) into y \[y=2^x\] Switch X and Y and take log of both sides. \[\log(x)=\log(2^y)\] Appy the rule of log and then divide both sides by 2 \[y=\log(x)/\log(2)\] Apply change of base rule \[y=\log_2(x)\] Now we convert it to 64° \[y=\log_2(64)\] 64 as 2^6 \[y=6\log_2(2)\] Evaluate the log_2(2) and we should get y=6. Our logarithmic function is \[y=\log_2(x)\] and the answer is when it is 64° is y=6.

NicktheBlade16:

Alright, thank you man, i appriacte it alot

Astrid1:

I hope I helped!!

Axolotlqueen:

my head hurts now